OpenStudy (anonymous):

2/3x+4=3/5x-2 they give me then answer already which is x=90 but they ask me to explain and I don't know how

OpenStudy (jonnyvonny):

The equation goes as this:\[\frac{ 2 }{ 3 }x + 4 = \frac{ 3 }{ 5 }x - 2 \] Is this correct?

OpenStudy (anonymous):

yes @JonnyVonny

OpenStudy (jonnyvonny):

Sorry, I was away for a while. Get the variables (the letters) alone by way of reverse PEMDAS. If you still need help, I'm here, otherwise, please inform me.

OpenStudy (anonymous):

ok thanks im good though I have figured it out @JonnyVonny